1. Identify and describe a process in your own organisation that can be used to generate relevant data.

2. Choose a quantitative variable associated with the process and discuss a relevant research question that could be investigated using the data.

3. Generate a data set for the quantitative variable comprising of at least 40 values.

4. Analyse the data set to include both relevant visual representations (graphs) and reduction methodologies (summary statistics)

5. State clearly the outcomes of your analysis.
Answer #1

These data will be used to construct a p control chart the subgroup size is n=100 the p values for each subgroup have been calculated and are shown in the table for in the 100 invoices inspected thus p=np/n=22/100=0.22 or 22% the p valus for the other days are calculates similarly

Cl=p

Ucl= p+3under root(pq/n)

Lcl= p-3under rrot(pq/n

P chart

Day Number

Invoices speed(n)

Number defective(np)

Fraction defective(p)

Q

1

100

22

0.22

0.78

2

100

23

0.23

0.77

3

100

25

0.25

0.75

4

100

36

0.36

0.64

5

100

25

0.25

0.75

6

100

24

0.24

0.76

7

100

45

0.45

0.55

8

100

21

0.21

0.79

9

100

25

0.25

0.75

10

100

28

0.28

0.72

11

100

26

0.26

0.74

12

100

21

0.21

0.79

13

100

23

0.23

0.77

14

100

26

0.26

0.76

15

100

36

0.36

0.64

16

100

35

0.35

0.65

17

100

38

0.38

0.62

18

100

39

0.39

0.61

19

100

21

0.21

0.79

20

100

23

0.23

0.77

21

100

32

0.32

0.68

22

100

32

0.32

0.68

23

100

30

0.30

0.70

24

100

23

0.23

0.77

25

100

23

0.23

0.77

26

100

25

0.25

0.75

27

100

21

0.21

079.

28

100

22

0.22

0.78

29

100

23

0.23

0.77

30

100

25

0.25

0.75

31

100

23

0.23

0.77

32

100

25

0.25

0.75

33

100

32

0.32

0.68

34

100

32

0.32

0.68

35

100

32

0.32

0.68

36

100

32

0.32

0.68

37

100

31

0.31

0.69

38

100

30

0.30

0.70

39

100

30

0.30

0.70

40

100

30

0.30

0.70